OpenStudy (anonymous):

find the square (4m-4)^2

OpenStudy (jadeishere):

(4m - 4)(4m - 4) 16m^2 - 16m - 16m + 16 16m^2 - 32m + 16

OpenStudy (anonymous):

Formula (a^2 - b^2) = a^2 - 2ab + b^2 & we have (4m - 4)^2 by comparing with this formula we will get =(4m)^2 - 2*(4m)*(4) + (4)^2 = 16m^2 - 32m + 16